OpenStudy (anonymous):

The computer system you need to input the function into only works in logarithms of base 10. Using complete sentences, explain how to convert your exponential function P(x) in a logarithmic one and then into a base 10 logarithm.

OpenStudy (anonymous):

P = 100 * 2^x log(P) = log(100 * 2^x) log(P) = log(100) + log(2^x) log(P) = log(100) + x log(2) log(P) = log(10^2) + x log(2) log(P) = 2 log(10) + x log(2) log(P) = 2 + x log(2) am i doing this right?

OpenStudy (anonymous):

@quinnisaghost

OpenStudy (anonymous):

maybe that'll help c~: